At the movie theatre, child admission is $6.60 and adult admission is 9.70On Monday, three times as many adult tickets as child tickets were sold, for a total sales of $1320.90How many child tickets were sold that day?
Found 2 solutions by ikleyn, MathTherapy:
Answer by ikleyn(52775)

.

x child tickets and 3x adult tickets were sold.


The total money equation is

    6.60x + 9.70*(3x) = 1320.90  dollars.


Simplify and find x

    6.60x + 29.10x = 1320.90

        35.70x     = 1320.90

             x     = 1320.90/35.70 = 37.


ANSWER.  37 child tickets.

Solved.



Answer by MathTherapy(10551)

At the movie theatre, child admission is $6.60 and adult admission is 9.70On Monday, three times as many adult tickets as child tickets were sold, for a total sales of $1320.90How many child tickets were sold that day? 

Let number of child tickets sold be C
Then number of adult tickets sold = 3C
We then get the following total-sales equation: 6.6C + 9.7(3C) = 1,320.9 
                                                  6.6C + 29.1C = 1,320.9
                                                         35.7C = 1,320.9
                           Number of child tickets sold, or